I need to know: How would i convert (x^2+y^2+2x-10=0) into (x-h)^2+(y-k)^2=r^2 format? I needed to find the radius and the center of the equation.
Answer by stanbon(75887)

convert (x^2+y^2+2x-10=0) into (x-h)^2+(y-k)^2=r^2 format
---------
You need to "complete the square" on the x terms and separately
on the y terms, as follows:
---------------
(x^2+2x+1) + (y^2) = 10+1
(x+1)^2 + (y-0)^2 = [sqrt(11)]^2
Center is at (-1,0)
Radius is sqrt(11)
Cheers,
Stan H.